# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v2

@@ A.3 Multi-scale Concat Construction and Training-Instance Generation
3: Initialize each molecule $m\in\mathcal{D}$ with a working graph $G_{m}$ where every atom is a node and $\mathrm{atom\_indices}(i)=\{i\}$
4: Compute initial statistics: for each $m$ and each edge $(u,v)\in E(G_{m})$, set $c\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(u)\cup\mathrm{atom\_indices}(v))$ and increment $\mathrm{stats}[c]$
@@
6:  Select $o_{t}\leftarrow\arg\max_{c}\;(\mathrm{stats}[c],\,c)$ (tie-break by lexicographic $c$ as in code)
@@
11:  Apply $o_{t}$ to all molecules that contain it: in each affected $G_{m}$, merge every edge $(u,v)$ whose union-fragment code equals $o_{t}$
12:  Update local pair statistics around merged nodes (remove old incident pairs; add new pairs)
13:  Set $\mathrm{stats}[o_{t}]\leftarrow 0$ (avoid re-selecting the same operation)
14:end for
Algorithm 2 Single-Atom Vocabulary Closure (SAVC) with AltForm backoff (implementation-level)
[…]
2:Output: basic single-atom token list $\Sigma_{\mathrm{atom}}$
@@
4:  Add X (standalone single atom) and X_AltForm (rare/atypical fallback)
5:  for each $v\in\mathrm{Val}(X)$ do
6:   for each attachment degree $d\in\{1,\dots,v\}$ do
7:    for each bond-type tuple $(b_{1},\dots,b_{d})\in\mathcal{B}^{d}$ do
8:     Construct a single-atom fragment SMILES with $d$ wildcard substituents (use branching if $d>1$), e.g. *X(*)(*)* with explicit bonds
9:     Add this motif_with_conn into $\Sigma_{\mathrm{atom}}$
@@
12:  end for
13:end for
14:Encoding-time backoff (MotifVocab.__getitem__): if a queried motif SMILES is not found after RDKit canonicalization, and it contains exactly one core atom (non-H, non-*), return the id of X_AltForm; otherwise return [UNK] (code uses id =1=-1 as unknown sentinel).
Algorithm 3 Materializing the connection-aware motif vocabulary (motif mining + SAVC integration)
1:Input: tokenizer corpus $\mathcal{D}$ (SMILES), operation prefix $\mathcal{O}_{\leq k}$, special tokens $\{\texttt{[PAD]},\texttt{[BOS]},\texttt{[EOS]},\texttt{[UNK]}\}$
2:Output: motif-pair list $\mathcal{V}=\{(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\}$ and token set $\Sigma$
3:for each SMILES $s\in\mathcal{D}$ do
4:  Parse $s$ to molecule $m$; build $G_{0}=(V_{0},E_{0})$ with atom/bond features
5:  Apply merges in order $\mathcal{O}_{\leq k}$ on a copy of $G_{0}$ to obtain motif graph $G=(V,E)$ where each node $v$ stores $\mathrm{atom\_indices}(v)\subseteq V_{0}$
6:  Assign each atom $i\in V_{0}$ to its motif node $\mathrm{bpe\_node}(i)\in V$
7:  for each original bond $(i,j)\in E_{0}$ do
8:   if $\mathrm{bpe\_node}(i)\neq\mathrm{bpe\_node}(j)$ then
9:    Add two connector nodes $i^{\star},j^{\star}$ with smarts=*; attach $i^{\star}$ to $i$ and $j^{\star}$ to $j$
10:    Set the bond types of $(i,i^{\star})$ and $(j,j^{\star})$ equal to the original bond type of $(i,j)$
11:    Include $i^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(i))$ and $j^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(j))$
@@
14:  for each motif node $v\in V$ do
15:   $v_{\mathrm{noConn}}\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(v)\cap V_{0})$
16:   $v_{\mathrm{withConn}}\leftarrow\mathrm{GraphSmiles}(G_{0}[\mathrm{atom\_indices}(v)])$ (contains * connectors)
17:   Add pair $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$ into a frequency counter
18:  end for
19:end for
20: Prepend special tokens; then integrate the SAVC basic vocabulary (Alg. 2) into the pair list
Algorithm 4 GraphBPE encoding with unknown-motif recovery (recursive split)
1:Input: SMILES $s$, operations $\mathcal{O}_{\leq k}$, motif vocabulary $\Sigma$ (lookup by canonical SMILES)
2:Output: ordered motif list $\big[(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\big]$
3:if $s$ contains “.then
4:  Split $s$ into fragments and encode each fragment independently; concatenate motif lists
5:end if
6: Build initial atom graph $G_{0}$ and initialize a binary merge tree at each atom node
7: Apply $\mathcal{O}_{\leq k}$ in order: for each code $o$, merge every edge whose union-fragment code equals $o$ (merge also updates the node’s merge tree)
8: Relabel nodes to integers and reorder them by scaffold-rooted BFS (Alg. 5)
9: Add connector * nodes for every cut bond to build connection-aware subgraphs (Alg. 3)
10: Collect motif tokens in BFS order; if a motif is not in $\Sigma$, mark it as unknown
11:while unknown motifs exist do
12:  Pop one unknown motif node and recursively split it into its two children in the stored merge tree
13:  For each child: if it is in $\Sigma$, emit it; else split further; stop at leaves (single-atom motifs), where SAVC backoff applies
@@
Algorithm 5 Scaffold-rooted BFS serialization (root selection + neighbor order)
1:Input: motif graph $G=(V,E)$ with node attribute $\mathrm{atom\_indices}(v)$
2:Output: a deterministic BFS ordering $\pi$ (used to relabel nodes)
3: Select root $r\leftarrow\arg\max_{v\in V}\big(|\mathrm{atom\_indices}(v)|,\;-\mathrm{id}(v)\big)$ (tie-break: smaller id)
4: Initialize queue with $r$, mark $r$ visited
5:while queue not empty do
6:  Pop front node $u$, append $u$ to ordering $\pi$
7:  Let $\mathcal{N}(u)$ be neighbors of $u$ not yet visited, sorted by ascending node id
8:  Push neighbors in $\mathcal{N}(u)$ into the queue and mark them visited
9:end while
10: Relabel nodes by $\pi$ so that the first visited node becomes id 0, the second becomes id 11, etc.
Algorithm 6 Multi-scale CamS concatenation and training-view generation
@@
2:Output: single-scale views $\{X_{s_{j}}\}_{j=1}^{S}$ and concatenated view $X_{\mathrm{concat}}$
3:for $j=1$ to $S$ do
4:  $X_{s_{j}}\leftarrow\iota_{s_{j}}.\mathrm{encode}(s).\mathrm{ids}$ (already includes [BOS] and [EOS])
5:end for
6: Build concatenated ids:
7:  $X_{\mathrm{concat}}\leftarrow[\texttt{bos\_id}]$
8:for $j=1$ to $S$ do
9:   Append the inner ids of $X_{s_{j}}$ (drop its BOS/EOS): $X_{s_{j}}[2{:}{-}1]$
10:  if $j<S$ then
11:    Append concat_id
@@
14:  Append eos_id
15:NTP loss masking: exclude [BOS], [EOS], [CONCAT] positions from prediction targets (implementation typically sets labels to -100 on these ids).
16:Padding/truncation: pad with [PAD] to max length; if truncating, keep the boundary tokens [BOS] and [EOS].
@@ Appendix B Theoretical Derivations
Overview. This section corresponds to the theoretical analysis in Sec. 3.3 and the supporting arguments referenced in Secs. 3.1 and 3.2.1. It supplements the main text with derivation-level details of (1) the DPI-based proof for the information loss induced by stochastic corruption (Proposition 3.1; Sec. B.1), (2) the direct supervision density / target-coverage counting under NTP vs. masked modeling in the multi-scale setting (Sec. B.2), and (3) the explicit structural-bias formulation that contrasts graph-transformer static bias with CamS causal constraint plus learned aggregation (Eq. (11); Sec. B.3).
@@
Setup.
For predicting token $x_{t}$, let $Z_{t}$ denote an unmasked evidence set (a collection of tokens) and let $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ be its stochastically masked version under a masking channel $\mathcal{M}$ (e.g., random node/substructure masking used in MLM/MNP). This yields a Markov chain
@@
[…]
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}),$ (15)
which proves Proposition 3.1. ∎
Remark (MLM limitations in NLP).
XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (Yang et al., 2019). Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (Clark et al., 2020).
Why this matters more on graphs (evidence-pattern uncertainty).
On graphs, the most predictive evidence for a token often lies in its local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings). Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance. This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (Li et al., 2023), (2) masking local subgraphs and predicting contextual properties (GROVER) (Rong et al., 2020), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (Liu et al., 2024). These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (Wettig et al., 2023).
CamS implication.
For CamS, coarse-scale token prediction under NTP conditions on an uncorrupted fine-to-coarse history by construction. Thus, relative to MNP which observes $\tilde{Z}_{t}$, CamS provides a higher-SNR conditioning set $Z_{t}$ that avoids stochastic evidence deletion. This does not claim that an arbitrary prefix always dominates the full bidirectional context; rather, it claims that masking as corruption strictly reduces the information of whatever evidence set is available, and CamS is designed so that the evidence set for coarse tokens is both uncorrupted and structurally aligned with the fine-grained neighborhood evidence.
B.2 Direct Supervision Density Analysis (Multi-Scale)
As a heuristic proxy for the richness of the learning signal, we compare the density of direct token-level supervision signals and how they accumulate over optimization.
Multi-scale notation.
Let the multi-scale CamS sequence consist of $S$ scales with lengths $\{T_{s}\}_{s=1}^{S}$ and total tokens $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$.
Where gradients originate.
@@ B.2 Direct Supervision Density Analysis (Multi-Scale)
Direct supervision density and accumulation.
@@
Equivalently, over $K$ optimization steps, the number of direct token-level targets scales as
$\mathcal{R}_{\text{NTP}}\propto K\cdot T_{\text{all}},\qquad\mathcal{R}_{\text{MNP}}\propto K\cdot\rho\,T_{\text{all}}.$ (17)
Thus, NTP provides about $1/\rho$ times more direct targets than MNP (e.g., $6.7\times$ when $\rho=0.15$).
@@
Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (Wettig et al., 2023). In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (Li et al., 2023). More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (Rong et al., 2020).
Why dense (and repeated) supervision can help.
Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (Clark et al., 2020). CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales). This reduces sensitivity to any single tokenization granularity and makes the effective supervision more robust.
@@ B.3 Structural Bias Formulation
We distinguish the structural injection mechanism in Eq. (11).
Graphformer (Hard Static Bias).
The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (Ying et al., 2021):
$\psi_{ij}^{\text{Graph}}=b_{\phi(d_{G}(i,j))}.$ (18)
@@
CamS uses a hard causal mask to enforce fine-to-coarse information flow:
$\psi_{ij}^{\text{CamS}}=\begin{cases}0&\text{if }(i\!\rightarrow\!j)\text{ is allowed by }\mathbf{M}_{\text{causal}}\\ -\infty&\text{otherwise.}\end{cases}$ (19)
Within the permitted region, connectivity is content-adaptive via $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$; hence coarse-scale tokens learn which fine-scale neighborhood details to aggregate, rather than uniformly mixing neighbors at a given distance.
Appendix C Training and Implementation Details
Overview. This section corresponds to the training and evaluation protocols in Secs. 3.2.1 and 4. It supplements the main text with implementation-level details of (1) the shared training framework and parallelization setup for pre-training vs. fine-tuning (Sec. C.1), (2) the complete pre-training configuration on Enamine675M (MultiScaler3375M) (Table 5), (3) the fixed fine-tuning setup used across MoleculeNet and MoleculeACE (Table 6), and (4) the unified hyperparameter search space (including strategy-specific grids) (Table 7).
C.1 Pre-training and Fine-tuning Hyperparameters
Implementation and framework.

### v3

@@ A.3 Multi-Scale Concatenation and Training Views
3: Initialize graph $G_{m}$ for each molecule $m\in\mathcal{D}$ (every atom is a node)
4: Compute initial pair statistics for all adjacent edges
@@
6:  Select operation $o_{t}\leftarrow\arg\max_{c}\;(\mathrm{stats}[c],\,c)$
@@
11:  Apply $o_{t}$ to all applicable edges in $\mathcal{D}$, merging nodes into larger motifs
12:  (Implementation Note: Merges are performed greedily. In case of overlaps, edges with smaller canonical atom indices are prioritized to ensure determinism.)
13:  Update local pair statistics around merged nodes
14:  Reset $\mathrm{stats}[o_{t}]\leftarrow 0$
15:end for
Algorithm 2 Single-Atom Vocabulary Closure (SAVC) & Back-off
1:Input: Element set $\mathcal{E}$, valences $\mathrm{Val}(X)$, bond types $\mathcal{B}$
2:Output: Basic single-atom vocabulary $\Sigma_{\mathrm{atom}}$
@@
4:  Add [X] (standalone) and [X_AltForm] (fallback) to $\Sigma_{\mathrm{atom}}$
5:  for each valence state and attachment pattern do
6:   Construct connection-aware SMILES with wildcards (e.g., *X(*)*)
7:   Add to $\Sigma_{\mathrm{atom}}$
8:  end for
9:end for
10:Encoding-Time Back-off Logic:
11:if queried motif is unknown AND contains exactly 1 core atom (element $X$) then
12:  return ID of [X_AltForm]
13:else
14:  return [UNK]
15:end if
Algorithm 3 Connection-Aware Vocabulary Materialization
1:Input: Corpus $\mathcal{D}$, operation prefix $\mathcal{O}_{\leq k}$, SAVC vocabulary $\Sigma_{\mathrm{atom}}$
2:Output: Full vocabulary $\Sigma$ (Map: Canonical SMILES $\to$ ID)
3:for each molecule $m\in\mathcal{D}$ do
4:  Apply $\mathcal{O}_{\leq k}$ to partition atoms into motif nodes $V$
5:  for each motif node $v\in V$ do
6:   Extract core fragment SMILES $v_{\mathrm{noConn}}$
7:   Extract connection-aware SMILES $v_{\mathrm{withConn}}$ (inserting * at cut bonds)
8:   Add $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$ to candidate set
9:  end for
10:end for
11: Filter candidates by frequency; Union with $\Sigma_{\mathrm{atom}}$ and Special Tokens to form $\Sigma$
Algorithm 4 Encoding with Recursive Unknown Recovery
1:Input: Molecule $M$, operations $\mathcal{O}_{\leq k}$, Vocabulary $\Sigma$
2:Output: Ordered token list
3:// Phase 1: Construction (Over-merge)
4: Initialize atom graph $G$. Assign each node a leaf BPETreeNode.
5:for operation $o\in\mathcal{O}_{\leq k}$ do
6:  Find all edges matching $o$.
7:  for edge $(u,v)$ in matches (ordered by node index) do
8:   Merge $u,v$ into $w$.
9:   Record merge history: $w.\text{tree}\leftarrow\text{Node}(\text{children}=[u.\text{tree},v.\text{tree}])$.
@@
12:// Phase 2: Serialization & Recovery
13: Serialize nodes $V_{s}$ via Scaffold-Rooted BFS (Alg. 5).
14: Initialize empty list $L$.
15:for node $v$ in BFS order do
16:  $L.\text{append}(\text{RecursiveResolve}(v,\Sigma))$
17:end for
18:return $L$
19:Function RecursiveResolve(node $v$, Vocabulary $\Sigma$):
20:if $v\in\Sigma$ then
21:  return [ID($v$)]
22:else
23:  // Unknown motif: backtrack using the BPE tree built in Phase 1
24:  Let $c_{1},c_{2}\leftarrow v.\text{tree}.\text{children}$
25:  return RecursiveResolve($c_{1},\Sigma$) + RecursiveResolve($c_{2},\Sigma$)
26:  // Base case: if single atom is unknown, apply SAVC back-off (Alg. 2)
27:end if
Algorithm 5 Scaffold-Rooted BFS (Intra-Scale Order)
1:Input: Motif Graph $G_{s}=(V_{s},E_{s})$
2:Output: Ordered list of nodes
3: Select root $r\leftarrow\arg\max_{v\in V_{s}}(\text{atom\_count}(v),-\text{node\_id}(v))$
4: Initialize Queue $Q\leftarrow[r]$, Visited $\leftarrow\{r\}$, Order $\leftarrow[]$
5:while $Q$ is not empty do
6:  $u\leftarrow Q.\text{pop\_front}()$
7:  Append $u$ to Order
8:  Get neighbors $\mathcal{N}(u)$, sort by node_id
9:  for $v\in\mathcal{N}(u)$ do
10:   if $v\notin\text{Visited}$ then
11:    $Q.\text{push\_back}(v)$, Visited.add($v$)
@@
15:return Order
Algorithm 6 Multi-Scale Concatenation (Inter-Scale Order)
1:Input: Molecule $M$, scales $\{s_{1},\dots,s_{M}\}$ (Fine $\to$ Coarse)
2:Output: Single-scale views $\{X^{(s_{j})}\}$, Multi-scale view $\mathbf{X}$
3:for $j=1$ to $M$ do
4:  $X^{(s_{j})}\leftarrow\text{Encode}(M,\mathcal{O}_{\leq s_{j}})$ (using Alg. 4)
5:end for
6:Construct Concatenated View:
@@
8:for $j=1$ to $M$ do
9:  Append tokens of $X^{(s_{j})}$ (excluding its own BOS/EOS)
10:  if $j<M$ then
11:   Append [CONCAT]
@@
14: Append [EOS]
15:return $\{X^{(s_{1})},\dots,X^{(s_{M})},\mathbf{X}\}$
@@ Appendix B Theoretical Derivations
Overview. This section corresponds to the theoretical analysis in Sec. 3.3. It supplements the main text with derivation-level details of: (1) The DPI-based proof for the information loss induced by stochastic corruption, including a discussion on Graph-specific evidence uncertainty (Sec. B.1); (2) The quantitative analysis of Direct Supervision Density, contrasting NTP vs. MNP and quantifying the gain from multi-view augmentation (Sec. B.2); (3) The explicit structural-bias formulation that contrasts graph-transformer static bias with CamS causal constraint plus learned aggregation (Sec. B.3).
@@
Setup and Markov Chain.
For predicting token $x_{t}$, let $Z_{t}$ denote an unmasked evidence set (e.g., the full causal history in CamS) and let $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ be its stochastically masked version under a masking channel $\mathcal{M}$ (e.g., random node/substructure masking used in MLM/MNP). This yields a Markov chain:
@@
By the Data Processing Inequality (DPI), for any Markov chain $X\!\leftrightarrow\!Z\!\rightarrow\!\tilde{Z}$, the data processing step $\tilde{Z}$ cannot increase the mutual information with the target $X$. Formally:
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}).$ (9)
This proves Proposition 3.1. ∎
Remark1: MLM limitations in NLP.
Our analysis aligns with classic critiques of Masked Language Modeling (MLM). XLNet (Yang et al., 2019) points out two intrinsic issues of corruption-based objectives: (1) the special token corruption (e.g., [MASK]) creates a pretrain-finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context. Relatedly, ELECTRA (Clark et al., 2020) shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions.
Remark 2: Graph-Specific Evidence Instability.
While the DPI inequality holds generally, the loss is particularly severe on graphs due to Evidence-Pattern Uncertainty. The most predictive evidence for a token often lies in its immediate local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings). Random masking $\mathcal{M}$ frequently removes not only the target token but also its critical neighborhood evidence (”co-masked”), making the conditional distribution $P(x_{t}|\tilde{Z}_{t})$ highly unstable across different masking patterns. This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability:
  • KPGT (Li et al., 2023): Masks structured units and adds global knowledge nodes to anchor the context.
  • GROVER (Rong et al., 2020): Masks local subgraphs and predicts contextual properties.
  • StructMAE (Liu et al., 2024): Uses structure-guided / curriculum masking rather than purely random masking.
These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (Wettig et al., 2023), but they still fundamentally operate on corrupted inputs. CamS Implication: In contrast, CamS conditions coarse-scale predictions on a structurally complete and uncorrupted fine-grained history. This ensures evidence stability and maximizes information gain without requiring auxiliary anchors.
B.2 Direct Supervision Density Analysis (Decomposition)
As a heuristic proxy for the richness of the learning signal, we compare the density of direct token-level supervision signals. To address the distinction between objective efficiency and data augmentation, we decompose the density gain into two factors: Intrinsic Efficiency and Systemic Multiplier.
Gradient Origins.
@@ B.2 Direct Supervision Density Analysis (Decomposition)
Factor 1: Intrinsic Objective Efficiency ($\times 1/\rho$).
@@
Thus, per pass, NTP provides about $1/\rho$ times more direct targets than MNP. For a typical masking rate $\rho=0.15$, this is a factor of $6.7\times$. This advantage is inherent to the NTP objective.
Factor 2: Systemic Augmentation Multiplier ($\times M$).
Crucially, the sequence representation of CamS naturally supports lightweight multi-view augmentation. We utilize 5 views per molecule (4 single-scale + 1 multi-scale) during pre-training. Let $\mathcal{R}$ be the total number of supervision targets provided by one molecule in one epoch:
$\mathcal{R}_{\text{MNP}}$ $\approx 1\times(\rho\cdot T_{\text{avg}})$ (11)
$\mathcal{R}_{\text{CamS}}$ $\approx 5\times(1\cdot T_{\text{avg}})$ (12)
Even against aggressive masking ($\rho=0.5$ as in KPGT), CamS provides $10\times$ more signals per molecule. While MNP can ostensibly adopt augmentation, re-encoding massive graphs for every view is computationally costlier than re-slicing sequences. CamS leverages the product of these factors, resulting in substantially higher sample efficiency.
@@
Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off (Wettig et al., 2023). KPGT reports that setting $\rho=0.5$ is only viable when global knowledge tokens help stabilize the conditioning context (Li et al., 2023). CamS bypasses this dilemma entirely, achieving maximum density ($\text{SD}\approx 1$) with zero corruption.
@@ B.3 Structural Bias Formulation
We distinguish the structural injection mechanism in Eq. (6) based on the bias term $\psi_{ij}$.
Graph Transformer (Hard Static Bias).
Standard Graph Transformers inject structure via static encodings, such as Shortest Path Distance (SPD), added to the attention scores (Ying et al., 2021):
$\psi_{ij}^{\text{Graph}}=b_{\text{SPD}(i,j)}.$ (13)
This imposes an isotropic prior: all atoms at distance $k$ are treated equally by the structural bias, regardless of their semantic content.
@@
CamS uses a hard causal mask to enforce fine-to-coarse information flow, as defined by the Inter-Scale Order:
$\psi_{ij}^{\text{CamS}}=\begin{cases}0&\text{if }i\in\text{Prefix}(j)\\ -\infty&\text{otherwise.}\end{cases}$ (14)
  • Bottom-up Composition: The Inter-Scale Order ensures that coarse motifs attend to their fine-grained constituents.
  • Anisotropic Learning: Within the permitted prefix ($\psi_{ij}=0$), connectivity is content-adaptive via $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$. This allows the model to dynamically select relevant fine-scale neighborhood details (e.g., focusing on a specific pharmacophore) rather than uniformly mixing neighbors, yielding a learned, anisotropic aggregation.
Appendix C Training Details and Benchmark Descriptions
Overview. This section supplements the experimental setup in Sec. 4. It provides: (1) Implementation details for the shared training framework (Sec. C.1); (2) Complete hyperparameter configurations for pre-training (Table 5) and fine-tuning (Tables 6 and 7); (3) Detailed descriptions of the downstream tasks in MoleculeNet and MoleculeACE (Sec. C.2).
C.1 Pre-training and Fine-tuning Implementation
Framework.